OpenStudy (anonymous):

how do I verify sin^2 (pi/2 -x) =tanx

OpenStudy (anonymous):

Can it be solved using a Ti84 Calculator?

OpenStudy (anonymous):

You can use a scientific calculator.

OpenStudy (anonymous):

check your statement.

OpenStudy (anonymous):

Thanks. yes, I typed the problem wrong. it should be sin^2(pi/2-x) +sin^2x=1

OpenStudy (anonymous):

\[2\sin ^2\left( \frac{ \pi }{ 2 }-x \right)=1-\cos 2\left( \frac{ \pi }{ 2 }-x \right)=1-\cos \left( \pi-2 x \right)=1+\cos 2 x\] \[=2 \cos ^2x,\sin ^2\left( \frac{ \pi }{ 2 }-x \right)=\cos ^2x\] solve further.